40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 255 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  4338   Tue Feb 22 14:37:24 2011 steveUpdateSAFETYLarisa received 40m safety training

Larisa Thorne received 40m lab specific, basic safety training. She will attend P. King's Basic Laser Safety Training Session tomorrow.

 

  4337   Tue Feb 22 11:53:38 2011 steveUpdateElectronicsRF Distribution box: assembly completed

Quote:

The mechanical assembly of RF distribution box is 99% complete.  Some of the components may be bolted to the teflon base plate if needed. 

All RF cables and DC voltage supply lines have been installed and tested.  I removed the terminal block which was acting as a distribution box for the common zero voltage line.  Instead I have used the threaded holes in the body of each voltage regulator.   This allows us to keep the supply lines twisted right up to the regulator and keeps the wiring neater.  The three regulator bodies have been wired together to provide a common zero potential point. 

I did a preliminary test to see if everything is functioning. All units are functioning well.  The output power levels may need to be adjusted by changing the attenuators. 

The 2x frequency multiplier outputs are not neat sine waves.  They seem to produce some harmonics, unlike the rest of the components.

I will post the measured power output at each point tomorrow.  The RF power meter could not be found in the 40m lab.  We suspect that it has found its way back to the PSL lab.

 

 http://www.timesmicrowave.com/wireless/index.shtml  

Frank is recommending these PhaseTrack-210 as phase stable low loss rf coax cables.

  4336   Tue Feb 22 00:41:34 2011 SureshUpdateElectronicsRF Distribution box: assembly completed

The mechanical assembly of RF distribution box is 99% complete.  Some of the components may be bolted to the teflon base plate if needed. 

All RF cables and DC voltage supply lines have been installed and tested.  I removed the terminal block which was acting as a distribution box for the common zero voltage line.  Instead I have used the threaded holes in the body of each voltage regulator.   This allows us to keep the supply lines twisted right up to the regulator and keeps the wiring neater.  The three regulator bodies have been wired together to provide a common zero potential point. 

I did a preliminary test to see if everything is functioning. All units are functioning well.  The output power levels may need to be adjusted by changing the attenuators. 

The 2x frequency multiplier outputs are not neat sine waves.  They seem to produce some harmonics, unlike the rest of the components.

I will post the measured power output at each point tomorrow.  The RF power meter could not be found in the 40m lab.  We suspect that it has found its way back to the PSL lab.

 

  4335   Tue Feb 22 00:18:47 2011 valeraConfiguration c1ioo and c1ass work and related fb crashes/restarts

I have been editing and reloading the c1ioo model last two days. I have restarted the frame builder several times. After one of the restarts on Sunday evening the fb started having problems which initially showed up as dtt reporting synchronization error. This morning Kiwamu and I tried to restart the fb again and it stopped working all together. We called Joe and he fixed the fb problem by fixing the time stamps (Joe will add details to describe the fix when he sees this elog).

The following changes were made to c1ioo model:

- The angular dither lockins were added for each optics to do the beam spot centering on MC mirrors. The MCL signal is demodulated digitally at 3 pitch and 3 yaw frequencies. (The MCL signal was reconnected to the first input of the ADC interface board).

- The outputs of the lockins go through the sensing matrix, DOF filters, and control matrix to the MC1,2,3 SUS-MC1(2,3)_ASCPIT(YAW) filter inputs where they sum with dither signals (CLOCK output of the oscillators).

- The MCL_TEST_FILT was removed

The arm cavity dither alignment (c1ass) status:

- The demodulated signals were minimized by moving the ETMX/ITMX optic biases and simultaneously keeping the arm buildup (TRX) high by using the BS and PZT2. The minimization of the TRX demodulated signals has not been successful for some reason.

- The next step is to close the servo loops REFL11I demodulated signals -> TMs and TRX demodulated signals -> combination of BS and PZTs.

The MC dither alignment (c1ioo) status:

- The demodulated signals were obtained and sensing matrix (MCs -> lockin outputs) was measured for pitch dof.

- The inversion of the matrix is in progress.

- The additional c1ass and c1ioo medm screens and up and down scripts are being made.

  4334   Mon Feb 21 23:00:06 2011 ZachSummaryelogrestarted

 again

  4333   Mon Feb 21 17:29:57 2011 ranaSummaryIOOMyterious data loss: FB needs investigation

Looks like there was a mysterious loss of data overnight; since there's nothing in the elog I assume that its some kind of terrorism. I'm going to call Rolf to see if he can come in and work all night to help diagnose the issue.

Untitled.png

  4332   Mon Feb 21 11:05:51 2011 ZachUpdateelogrestarted

 I restarted the elog using the script.

  4331   Sun Feb 20 21:22:33 2011 rana, kiwamu, valeraConfigurationIOOMC Servo Change

For some reason, Kiwamu forced us to change the MC servo electronics today. We are now combining it with the FSS box.

The MC Servo by itself was locking by just driving the NPRO PZT. Becuase of the ~30 kHz mechanical resonances of that system, our badnwidth is limited. To get higher bandwidth, we can either use a wideband frequency shifter like the AOM or just use the ole FSS combo of PZT/EOM. The old MC servo was able to get 100 kHz because it used the AOM.

So we decided to try going through the FSS box. The MC servo board's FAST output now goes into the IN1 port (500 Ohm input impedance) of the TTFSS box. This allows us to use the FSS as a kind of crossover network driving the PZT/EOM combo.

At first it didn't work because of the 5V offset that Jenne, Larisa, Koji, and Suresh put into there, so I cut the wire on the board that connected the power to the summing resistor and re-installed the MC Servo board.

We also removed the old Jenne-SURF 3.7 MHz LP between the MC mixer and servo. Also removed the Kevin-box (1.6:40) stuck onto the NPRO PZT.

We have yet to measure the UGF, but it seems OK. The PCDRIVE is too high (~5-6V) so there is still some high frequency oscillation. Needs some investigation.

* To get the FSS SLOW servo to work (change NPRO temperature to minimize PZT drive onto NPRO) I set the setpoint to 5V in the script so that we operate the FSS box output at 5V mean. I set the threshold channel to point to MC_TRANS_SUM instead of RC_TRANSPD. I also had to fix the crontab on op340m so that it would point to the right scripto_cron script which runs the FSSSlowServo, RCThermalPID.pl, etc. I also had to fix scripto_cron itself since it had the old path definitions and was not loading up the EpicsTools.pm library.

** Also, I was flabbergasted by the dog clamping on the last turning mirror into the MC. Barely touching the mount changes the alignment.

  4330   Sat Feb 19 05:25:20 2011 SureshUpdateElectronicsRF: Distribution box

Most of the RF cables required for the box are done.   There are two remaining and we will attend to these tonight. 

We expect to have finished the mechanical assembly by Sunday and start a quality test on Monday.

 

 

  4329   Sat Feb 19 01:58:20 2011 ranaSummaryElectronicsLow Noise BJT Pre-Amp

Frank put his low noise preamp info here.

I suggest that we build these (using Altium) but replace the cheapo transistors with the high class MAT03 matched BJT pair from Analog Devices.

This will allow us to have a pre-amp better matched to the noise of the mixers down to low frequency.

  4328   Fri Feb 18 20:17:07 2011 JoonhoSummaryElectronicsIsolation of Voltage regulator

Today I was working on RF distribution box.

So far I almost finished to electronically isolate voltage regulators from the box wall by inserting mica sheet, sleeve, and washers.

 

The problem I found is the resistance between wall and the voltage regulator is order of M ohms

I checked my isolation (mica sheet and sleeve and washer) but there is no problem there.

But I found that the power switch is not completely isolated from the wall.( around 800 kohm)

and that the resistance between the regulator and the wall is smaller for the regulator closer to the power switch

and greater for the regulator less closer to it.

So I think we need to put washer or sleeve to isolate the powersitch electronically from the box wall.

Suresh or I will fix this problem

[ To Suresh, I can finish the isolation when I come tomorrow. Or you can proceed to finish isolation.]

  4327   Fri Feb 18 20:06:59 2011 kiwamuSummarySUScheck f2p function on ETMX

 The plot below shows how the f2p filters work.

At -2 min I turned on the f2p filters.

 f2p_ETMX.png

  4326   Fri Feb 18 18:46:08 2011 kiwamuSummarySUSf2p done on ETMX and ITMX

The f2p measurements are done on ETMX and ITMX with the real time lockin systems.

I don't explain what is the f2p measurement in this entry, but people who are interested in it can find some details on an old elog entry here or somewhere on DCC.

So far the resultant filters looked reasonable compared with the previous SRM f2p filters.

 

- backgrounds -

  Some times ago I found that the coils on ETMX had not  been nicely balanced, and it made a POS to angle coupling when I tried green locking (see here).

In addition to that, accuracy of A2L kind of measurement including the dithering techniques depend on how well the coils are balanced.  Therefore we need to balance the coils basically at all the suspended optics.

There used to be a script for this particular purpose, called f2praio.sh. This script does measure the imbalances and then balance the coils.

However this time I used the realtime lockin system to measure the imbalances instead of using the old f2p script.

One of the reasons using the real time system is that,  some of the ezca and tds commands for the old script don't work for some reasons.

Therefore we decided to move on to the real time system like Yuta did for the A2L measurement a couple of months ago.

The f2p measurement finally gives us parameters to generate a proper set of filters for POS and also the coil gains. We apply those filters and the gains in order to eliminate the POS to angle coupling and to balance the coils.

 

- results -

The followers are the resultant filters and coil gains.

The plots below show new f2p filters according to the measurement.

f2p_ITMX.png       f2p_ETMX.png

 

ITMX (assuming pendulum POS has f0 = 1 Hz and Q = 1)

ULPOS  fz = 1.009612   Qz = 1.009612 

URPOS fz = 1.125965   Qz = 1.125965  

LLPOS  fz = 0.873725   Qz = 0.873725    

LRPOS  fz = 0.974418   Qz = 0.974418  

C1:SUS-ITMY_ULCOIL_GAIN      -1.103044

C1:SUS-ITMY_URCOIL_GAIN      0.884970

C1:SUS-ITMY_LLCOIL_GAIN      0.950650

C1:SUS-ITMY_LRCOIL_GAIN      -1.060326

 

 

ETMX (assuming pendulum POS has f0 = 1 Hz and Q = 1)

ULPOS  fz = 1.055445   Qz = 1.055445   

URPOS  fz = 1.052735   Qz = 1.052735   

LLPOS  fz = 0.944023   Qz = 0.944023   

LRPOS  fz = 0.941600   Qz = 0.941600   

C1:SUS-ETMX_ULCOIL_GAIN      -0.887550

C1:SUS-ETMX_URCOIL_GAIN      1.106585

C1:SUS-ETMX_LLCOIL_GAIN = 1.07233

C1:SUS-ETMX_LRCOIL_GAIN      -0.931013

  

The precision of the coil gains looked something like 1% because every time I ran the measurement script, the measured imbalances fluctuated at this level.

The precision of the filter gain at DC (0.01 Hz) could be worse, because the integration cycles for the measurement are fewer than that of the coil gains done at high frequency (8.5 Hz).

Of course we can make the precisions by increasing the integration cycles and the excitation amplitudes, if we want to.

  4325   Fri Feb 18 17:52:25 2011 josephb, valeraUpdateCDSc1ass updated

We updated the c1ass model to include the BS.  We removed the dither excitation of the PZTs.  PZT control goes to epics. To do this, modified the /cvs/cds/caltech/target/c1iscaux/PZT_AI.db file.  We basically have it sum both the existing epics slider and our new output from c1ass.

More importantly we updated the color scheme.

We compiled and tested the Dolphin and RFM which work.

I should note we can't figure out why testpoints are not working properly with just this model.  Alex and Joe spent well over an hour trying to debug it to no success.  Current workaround is to add what channels you want from c1ass to the DAQ recording.  Other testpoints on other models appear to be working.

Attachment 1: c1ass_updated.png
c1ass_updated.png
  4324   Fri Feb 18 15:05:49 2011 kiwamuUpdateGreen Lockingtransfer function of angle to beat note (length)

[Koji and Kiwamu]

 We took transfer functions (TF) from the angle excitations at ETMX and ITMX to the green beat note signal (i.e. angle to length TF).

It turned out that the coupling from ETMX_PIT is quite large.

I wonder how f2p of the ETMX changes this coupling. We'll see.

 

a2l_TF.png

 

The plot above shows a set of the transfer functions from the angle excitation to the green beat note.

Note that the y-axis has not been calibrated, it is just a unit of counts/counts.

You can see that the TF from ETMX_PIT to the beat (red cruve) is larger than the others by about a factor of 10 over most of the frequency range.

This means that any PIT motions on ETMX can be coupled into the green beat signal somewhat over the wide frequency range.

It looks having a resonance at 1.5 Hz, but we don't exactly know why.

At that time the coil gains on only ITMX were tuned by applying f2p filters, but ETMX wasn't because of a technical reason coming from epics.

 

- - - - measurement conditions

  * PSL laser was locked to X arm by feeding back the IR PDH signal to MC2.

  * the green laser was locked to Xarm as usual.

  * took the green beat note signal (approximately 0 dBm) into Rana's MFD with the cable length of about  6 m.

  * the output from the MFD was connected to XARM_COARSE channel without a whitening filter.

  * excitation signal was injected into either ASC_PIT or ASC_YAW. The excitation was Gaussian noise with frequency band of 10 Hz and amplitude of 300 counts.

  * only ITMY had the f2p filters, which balance the coil gains all over the frequency.

 

  4323   Fri Feb 18 13:41:22 2011 josephbUpdateCDSCDS fixes

I talked to Alex today and had two things fixed:

First the maximum length of filter names (in the foton C1SYS.txt files in /chans) has been increased to 40, from 20.  This does not increase EPICS channel name length (which is longer than 20 anyways).

This should prevent running into the case where the model doesn't complain when compiled, but we can't load filters.

Additionally, we modified the feCodeGen.pl script in /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/util/ to correctly generate names for filters in all cases.  There was a problem where the C1 was being left off the file name when in the simulink .mdl file the filter was located in a box which had "top_names"  set.

  4322   Fri Feb 18 00:35:21 2011 kiwamuSummaryGeneralto do list for Feb. 18th

DAYTIME

  -  more precise F2P measurement and modify lockin simlink model (Kiwamu)

  -  run  C1ASS to check it (Valera)

  -  take care of CDS (Joe)

  -  MC mode matching (Jenne/Koji)

   -  RF stuff  (Suresh)

  -  mode matching for doubling crystal at PSL table (low priority)

  -  OPLEV (low priority)

NIGHITTIME

 -  update the noise spectra of green locking

 -  make noise budgets

  4321   Fri Feb 18 00:13:55 2011 kiwamuUpdateCDSRe:Daqd was rebuilt, now reverted.

THANK YOU, JOE !!! 

Quote:

As one of the trouble shooting steps for the daqd (i.e. framebuilder) I rebuilt the daqd executable.

  4320   Thu Feb 17 23:56:53 2011 josephbUpdateCDSDaqd was rebuilt, now reverted.

As one of the trouble shooting steps for the daqd (i.e. framebuilder) I rebuilt the daqd executable.  My guess is somewhere in the build code is some kind of GPS offset to make the time correct due to our lack of IRIG-B signal.

The actual daqdrc file was left untouched when I did the new install, so the symmetricom gps offset is still the same, which confuses me.

I'll take a look at the SVN diffs tomorrow to see what changed in that code that could cause a 300000000 or so offset to the GPS time.

 

 

  4319   Thu Feb 17 23:41:46 2011 ranaFrogsDAQFrames Directory got the wrong name: Data unreachable

DTT stopped working for recent data. An 'ls' in the frames/full/ directory reveals:

drwxr-xr-x 2 controls controls 258048 Feb  3 12:26 9807
drwxr-xr-x 2 controls controls 258048 Feb  4 16:13 9808
drwxr-xr-x 2 controls controls 262144 Feb  5 19:59 9809
drwxr-xr-x 2 controls controls 258048 Feb  6 23:46 9810
drwxr-xr-x 2 controls controls 258048 Feb  8 03:33 9811
drwxr-xr-x 2 controls controls 262144 Feb  9 07:19 9812
drwxr-xr-x 2 controls controls 253952 Feb 10 11:06 9813
drwxr-xr-x 2 controls controls 266240 Feb 11 14:53 9814
drwxr-xr-x 2 controls controls 266240 Feb 12 18:39 9815
drwxr-xr-x 2 controls controls 266240 Feb 13 22:26 9816
drwxr-xr-x 2 controls controls 262144 Feb 15 02:13 9817
drwxr-xr-x 2 controls controls 253952 Feb 16 05:59 9818
drwxr-xr-x 2 controls controls 241664 Feb 17 09:46 9819
drwxr-xr-x 2 controls controls  28672 Feb 17 12:22 9820
drwxr-xr-x 2 controls controls  32768 Feb 17 15:06 6663
drwxr-xr-x 2 controls controls  73728 Feb 17 23:39 6664
controls@fb /frames/full $ date
Thu Feb 17 23:39:27 PST 2011

  4318   Thu Feb 17 23:11:40 2011 ranaUpdateElectronicsVCO Frequency Noise Measurement with the MFD

This is the 140 ft. MFD measurement of the VCO phase noise. It is open loop and so should be a good measurement. The RMS is 30 Hz integrated down to 2 mHz.

I don't know why this doesn't agree with Suresh's measurements of the same thing which uses the PLL feedback method.

In BLUE, I also plot the frequency noise measured by using a Stanford DS345 30 MHz func. generator. I think that this is actually the noise of the FD (i.e. the SR560 preamp) and not the DS345. Mainly, it just tells you that the PINK VCO noise measurement is a real measurement.

I calibrated it by putting in a 5 kHz_pp triangle wave on the sweep of the DS345 and counting the counts in DV.

Attachment 1: vco.png
vco.png
  4317   Thu Feb 17 22:51:04 2011 josephb, valeraSummary dither alignment model

We made a model for the dither angular stabilization system c1ass.mdl. The attached file shows the diagram.

The idea is to dither a combination of 6 optics (ETMs, ITMs, PZTs) at different frequencies and demodulate three PDs (TRX, TRY, REFL11I). Then form the DOFs from demodulted signals, filter, and send each DOF to a combination of optics.

This is enough to get started with arm cavities alignment (we may need to add the BS for the Y arm). More optics and PD can be added as they become available and/or needed.

The DAC for the fast PZT  are not connected and have to be commissioned.

Attachment 1: ass-model.png
ass-model.png
  4316   Thu Feb 17 14:52:27 2011 JenneUpdateIOOMC alignment work

I worked a little bit more on optimizing the mode matching to the MC, but it's still not great.  I've only gotten a visibility of ~45%, but Koji said that it used to be ~87%.  So there is a long way to go.  Kiwamu said he can work with the lower-power configuration for a few days, and so my next step will be to measure the beam profile (stick a window in the path, and look at the refl from the window....that way we don't get thermal lensing from transmission through an optic), and redo the mode matching calculation, to figure out where the last lens should actually sit.

Quote:

So.... Kiwamu and I were concerned (still a little concerned) that ETMY is not damping as nicely as it should be.  (It's fine, but the UL rms is ~5, rather than ~1 or less. BURT restores by Kiwamu didn't change anything.) Anyhow, I was heading out to push the annoying ribbon cables more firmly into the satellite adapter board things that are tied to the racks in various places (The back of 1X5 for the corner optics and the end station racks for the ETMs).  The point was to push in the ETMY one, but while I was out in the lab and thinking about it, I also gave all of the corner connectors (MC1, MC2, MC3, ITMx, ITMY, BS, PRM, SRM) a firm push. 

Kiwamu noticed that when I did this, the Mode Cleaner alignment got a little bit worse, as if the connection to the satellite adapter boards hadn't been great, I pushed the connectors in and the connection got better, but we also got a bit of a DC offset in the MC alignment.  Anyhow, the MC_TRANS power went down by ~2, to about the place it had been before Kiwamu adjusted the position of the lens in between the zigzag mirrors.  (I don't know if Kiwamu elogged it earlier, but he scooted the lens a teensy bit closer in the optical path to the Mode Cleaner). 

To counteract this loss in MC transmitted power as a result of my connector actions, I went back to the PSL table and fiddled with the zigzag steering mirrors that steer the beam from the PSL table over to the mode cleaner.  I got it a little better, but it's still not perfect.

Kiwamu has noted that to improve the mode matching into the Mode Cleaner with the new PMC in place, we might have to move the lens which is currently between the zigzag steering mirrors, and put it after the second mirror (so in between the last steering mirror and the pickoff window that sends a piece of the beam over to PSL_POS and PSL_ANG).  This will make the waist between MC1 and MC3 tighter. 

Moral of the story:  To improve IMC mode matching we need to move the last lens closer in the optical path to the mode cleaner waist. Twiddle with zigzag steering mirrors to optimize.

 

  4315   Thu Feb 17 14:17:27 2011 SureshUpdateElectronicsRF Distribution box and REFL11

 

The Distribution box is several steps nearer to completion.

 

1) Soldered capacitors and DC power lines for four units of the distribution box.

2) mounted all the components in their respective places.

3) Tomorrow we prepare the RF cables and that is the last step of the mechanical assembly. 

4) we plan to test both the generator and distributon parts together.

 

 

 

REFL 11

 

[Kevin, Suresh]

Kevin took a transfer function of the newly assembled PD and noticed that the frequency has shifted to 14.99  freom 11. MHz.

We needed to find the current RLC combination.  So we  removed the ferrite core from L5 rendiring it to its aircore value of  0.96/muH. We then used this to find the Capacitance of the PD (117pF)

We  used this value to compute the inductance required to achieve 11.065MHz  which turned out to be 1.75microH.

This was not reachable with the current L5 which is of the type  143-20J12L (nominal H=1.4 micro Henry).

We therefore changed the inductor to SLOT 10 -3-03. It is a ferrite core, shielded inductor with a plasitc sleeve. Its nomial valie is 1.75 microH

We then tested the DC output to see if here is a response to light. There was nonel. l

The problem was traced to the new inductor.  Surprisingly the inductor coil had lost contact with the pins.

I then replacd the inductor and checked again.  The elecronics seems to work okay..   but there is a very small signal 0.8mV for 500microW. 

There seems to be still something wrong with the PD or its electronics.

 

 

  4314   Thu Feb 17 13:20:06 2011 SureshUpdateVIDEOSome more labels

[Larisa, Kiwamu, Steve and Suresh]

 

  We continued the labeling of video cables. All exiting cables which are going to be used used in the new scheme have been labeled.

We also labeled the cables running from the video mux to the TV monitors in the computer room. Some of these will be removed or reallocated.

We will continue next Wednesday (after the meeting) and will lay cables that are most urgently required. 

 i

  4313   Thu Feb 17 11:51:14 2011 josephbUpdateCDSLockin filter names too long - broke loading

Problem:

Could not load filters into the C1:SUS-ETMX_LOCKIN1_SIG filter bank.

Reason:

Apparently the filter bank name was too long.  I'm not sure why this isn't caught by the real time code generator, I'm planning on asking Alex and Rolf about it today.

Solution:

Reduce the name of the components.  Basically LOCKIN1 needs to become something like LOCK1 or LIN1.

 

In related news, it looks like the initial filters are hard coded to be 2048 Hz.  Given that they start out empty they won't cause things to break immediately, and if you're editing the file you can update the rate as you add the filter.  I'll also bring this up with Alex and Rolf and see if the RCG can't be more intelligent about its filter generation.

 

  4312   Thu Feb 17 11:49:48 2011 josephbUpdateCDSFront end start/stop scripts go to /scripts/FE again

I modified the core/advLigoRTS/Makefile to once again place the startc1SYS and killc1SYS scripts in the scripts/FE/ directory.

It had been reverted in the SVN update.

  4311   Thu Feb 17 11:20:04 2011 josephbUpdateCDSstart scripts no longer need sudo

I've modified the rc.local file to run the IOC codes as controls, which means they no longer write root permission log files on startup.

The awgtpman, which was the other permission issue with the start scripts, is started by a run script now.  This new version seems to be content to keep the permissions of the current log file, which is set to controls.

This should prevent the issue of sudo wiping your path environment variable for just that command. (Try "sudo which burtwb" versus "which burtwb" for example).  This apparently a security feature of sudo.

If you should happen to use sudo to run a start script, the easiest solution to fix the permissions is just got to the target directory (type "target") and run "sudo chown controls:controls -R *" on one of the workstations (the front ends don't handle the groups properly at the moment).

This should allow the scripts to properly use burtrb and burtwb to write and backup burt files.

  4310   Thu Feb 17 00:43:13 2011 kiwamuSummaryGeneralto do list for tomorrow

Let's share our tasks !

 

==== daytime ===

 . mode matching for MC  (Jenne/Koji)

 . mode matching for doubling crystal on PSL table (Suresh/Koji)

 . f2p adjustment (Kiwamu)

 . fix daq and CDS issues (Joe)

 . increase oplev gain (low priority)

 . make ITMY camera nicer (Steve)

 . c1ass simlink model (Valera/Joe)

 . Bounce Roll notches (Suresh)

 

==== nighttime (after 7pm) ====

 . align everything   (at first green beam, then X arm cavity and finally IR beam)

 . update the noise spectrum of the green locking

 . estimate the noise from angle to length coupling

  4309   Wed Feb 16 23:54:34 2011 ranaUpdateCDSToday's CDS problems
  1. ETMX cannot load his filter coefficients. Even if I change the file, the load button doesn't work. I tried changing the lockin filters but they won't load.
  2. The ETMX filter modules appear to have 2048 for some of the modules and 16384 for some of the others. How come the make script doesn't make them all 16384?
  3. There are a bunch of kill/start scripts in the scripts/ directory instead of in the scripts/FE/ directory. Did this get reverted after a new code exchange?
  4. I tried restarting the code using c1startscx, but that doesn't work correctly. It cannot find the burtrb and burtwb files even though they are in the normal path.
  5. Kiwamu was using a bunch of cockamamie filters I found.
  6. I can't get any minute trend data. I tried on rossa, rosalba, and op440m.
MC damp dataviewer diaggui AWG c1lsc c1ioo c1sus c1iscex c1iscey RFM The Dolphins Sim.Plant Frame builder TDS Cabling
                             
  4308   Wed Feb 16 12:16:14 2011 josephbUpdateCDSFixed Optical level SUM channel names

[Joe,Valera]

Valera pointed out the OPLEV SUM channels were incorrect.  We changing the optical level sum channel to _OPLEV_SUM when it should have been OL_SUM.  This has been fixed in the activateDAQ.py script.

  4307   Wed Feb 16 10:35:40 2011 Larisa ThorneUpdateIOOWFS quantum efficiency as a function of angle

 Here is the followup on Jenne's February 14th, 2011 update on the quantum efficiency measurements of WFS2.

http://nodus.ligo.caltech.edu:8080/40m/4289

 

Attached is a PDF of my calculations, based on measurements ranging between 0-25 degrees in 5 degree increments.

The graph at the bottom plots these angles versus the calculated quantum efficiency at each point and the responsivity. Since quantum efficiency and responsivity only differ by a factor of some natural constants (lamda, e, h, c), I used a graph with two vertical axes, because the points would be plotted at essentially the same location if quantum efficiency (%)  and responsivity (Amps/Watts) were graphed on two separate plots.

The calculated values for quantum efficiency based on my measurements (labelled "ExpAverage") were pretty close to what Jenne had calculated in earlier attempts, which was around 60%. Just to test, I compared my quantum efficiency result against the calculation of quantum efficiency using the responsivity value for silicon, 0.5 Amps/Watt, which is labelled as "Spec". Comparison of "ExpAverage" and "Spec" shows that they differ by only about 2%, so I conclude that the theoretical quantum efficiency calculated using a given responsivity agrees with my measurement-based experimental result.

Attachment 1: QEcalcs.pdf
QEcalcs.pdf
  4306   Wed Feb 16 02:04:11 2011 kiwamuUpdateASCIR beam alignment

[Jenne and Kiwamu]

 This time we aligned the vertical angle (not the translation) of the IR beam so that the transmitted light from BS shoots the center of ETMY.

The idea is to use ETMY as a beam pointing reference instead using IP_ANG, assuming the translation is not so bad.

As a result it looks like we are wining. A quick A2L test on ITMX_PITCH showed a small off-centering at sub-milimeter level.

 

 We are concluding that the initial beam after PZT2 had been pointing downward somehow.

Before doing this whole job, we checked the spot shape on IP_POS to see if the beam is clipped or not. It was a round shape, which means no clipping around MMT.

But on the other hand, the spot on IP_ANG had been clipped more than half of its bottom as Suresh reported on his elog (see here).

I found that this clipping is able to be fixed by moving the beam angle upward. I guess the clipping happened at one of the steering mirror in the ETMY chamber.

According to these information, we imagined that the beam was somehow pointing downward after PZT2.

So we started aligning the beam by touching only PZT2 for vertical direction. Then we found a beam spot on ETMY's suspension frame, and brought it to the center.

Then we aligned BS and X arm for this new beam axis. The it resulted a small off-centering on pitch.

Once the MC fully gets back, we will examine the TRX degradation with this configuration.

  4305   Wed Feb 16 01:03:59 2011 JenneUpdateIOOMC alignment work

So.... Kiwamu and I were concerned (still a little concerned) that ETMY is not damping as nicely as it should be.  (It's fine, but the UL rms is ~5, rather than ~1 or less. BURT restores by Kiwamu didn't change anything.) Anyhow, I was heading out to push the annoying ribbon cables more firmly into the satellite adapter board things that are tied to the racks in various places (The back of 1X5 for the corner optics and the end station racks for the ETMs).  The point was to push in the ETMY one, but while I was out in the lab and thinking about it, I also gave all of the corner connectors (MC1, MC2, MC3, ITMx, ITMY, BS, PRM, SRM) a firm push. 

Kiwamu noticed that when I did this, the Mode Cleaner alignment got a little bit worse, as if the connection to the satellite adapter boards hadn't been great, I pushed the connectors in and the connection got better, but we also got a bit of a DC offset in the MC alignment.  Anyhow, the MC_TRANS power went down by ~2, to about the place it had been before Kiwamu adjusted the position of the lens in between the zigzag mirrors.  (I don't know if Kiwamu elogged it earlier, but he scooted the lens a teensy bit closer in the optical path to the Mode Cleaner). 

To counteract this loss in MC transmitted power as a result of my connector actions, I went back to the PSL table and fiddled with the zigzag steering mirrors that steer the beam from the PSL table over to the mode cleaner.  I got it a little better, but it's still not perfect.

Kiwamu has noted that to improve the mode matching into the Mode Cleaner with the new PMC in place, we might have to move the lens which is currently between the zigzag steering mirrors, and put it after the second mirror (so in between the last steering mirror and the pickoff window that sends a piece of the beam over to PSL_POS and PSL_ANG).  This will make the waist between MC1 and MC3 tighter. 

Moral of the story:  To improve IMC mode matching we need to move the last lens closer in the optical path to the mode cleaner waist. Twiddle with zigzag steering mirrors to optimize.

  4304   Tue Feb 15 21:45:08 2011 ranaUpdateIOOThe MC TRANS Story

I forgot to elog that last night I touched up the MC2_TRANS QPD setup. I was perplexed by it always going out of alignment so I investigated.

I found that the fork clamp for the steering mirror for the QPD was not tightened. Shame. The beam diameter was equal to the aperture of the QPD and was clipping. Double shame.

I added a lens and tightened the mounts and centered the beam at ~9 PM yesterday. You can see in the attached trend that the measured power went up by ~10%.

Later, there's a big gap where Valera and Steve change out the PMC. You can see that the MC REFL voltage goes from 4.5 V to 5 V (10% increase in the power delivered to the MC).

There's essentially no change in the total transmission - this indicates that although the PMC transmission is now higher by ~10%, the matching to the IMC has been degraded by an equivalent fraction.

Needs some mode matching work.

Attachment 1: a.png
a.png
  4303   Tue Feb 15 17:48:22 2011 JenneUpdateIOOMode Cleaner resonating again

[Valera, Jenne]

After Steve and Valera switched out the PMC, the Mode Cleaner resonance needed to be brought back.  We spent some time playing with the 2 steering mirrors directly after the PMC, to get the beam through the EOM, and to achieve flashing in the MC.  Valera then adjusted those 2 steering mirrors to minimize MC_REFL_DC.  I did a little bit more, and it's kind of close now, but we're only at ~half normal transmitted power.  Since the 2 steering mirrors after the PMC are so close together, the beam alignment is pretty sensitive to even small touches.  So it's probably time to move on to using the last zigzag steering mirrors on the PSL table, since they're farther apart. 

I have to head out for a little while, but I'll be back in a few hours. Kiwamu said he might continue the alignment into the MC, if he needs the IFO.  Also, we should measure the power before and after the EOM, just to confirm that we're getting through it optimally.  The beam looks good after the EOM, and the MC is resonating, so it should be fine, but it can't hurt to check.

  4302   Tue Feb 15 15:06:25 2011 josephbUpdateCDSCDS todo list for tomorrow morning

Currently, there is a test directory called /opt/rtcds/caltech/c1/new_core where we have the latest svn checkout.  Tomorrow (after everything works), it will become the core directory.

1) Modify on the fb machine the /diskless/root/etc/ld.so.cache file.  This is done by logging into fb, going to /etc/ld.so.conf.d/, modifying epics-x86_64.conf to only have .10 stuff , and running sudo /sbin/ldconfig.  Copy the newly generated /etc/ld.so.cache file to /diskless/root/etc/.

2) Modify the rc.local file on the fb machine in /diskless/root/etc/ to take advantage of the new subscripts and init.d/ start scripts.

3) Add the no_rfm_dma to all the iop models (c1x01,c1x02,c1x03,c1x04,c1x05).

4) Rebuild all front end models with new code.  Install.

5) Build awgtpman and mx_streams with new code.

6) Rerun activateDaq.py (to fix channel names from all the rebuilt code).

7) Double check Burt request files have the switch fix.

8) Restart the front ends.

9)Restart the frame builder.

9) Check channels, exitations, RFM connections.

10) Check Monit is working.

  4301   Tue Feb 15 11:57:06 2011 steve, valeraConfigurationPSLPMC swap

 We swapped the PMC s/n 2677 for s/n lho006.

The table below summarizes the power levels before and after the PMC swap.

  old new
Ptrans 1.32 W 1.42 W
Transmission 85 % 91.5 %
Refl PDDC locked/unlocked 5.0 %  4.3 %
Loss 7-8 % 2-3 %
Leakage out of the back 10 mW 0.3 mW

 

- The power into the PMC (1.67 W) was measured with Scietech bolometer before the first steering PMC mirror. The leakage through the steering mirrors was measured with Ophir power meter to be 12+8 mW. There is also a lens between the mirrors which was not measured. 

- The power through the PMC was measured after the doubler pick off (105 mW), steering mirror (4 mW), and lens (not measured).

- The estimated reflection from four lens surfaces is 1-2% hence 1% uncertainty in the losses in the table.

- The beams into the PMC and on REFL PD were realigned. The beams downstream of the PMC are blocked as we did not realigned the PMC and doubler paths.

- The trans PD ND filters were removed. The VDC=1.28 V now.

- The NPRO current is 2.102 A

 

Atm 1 old

Atm2  new

Attachment 1: P1070421.JPG
P1070421.JPG
Attachment 2: P1070423.JPG
P1070423.JPG
  4300   Tue Feb 15 11:56:17 2011 josephbUpdateCDSUpdated some DAQ channel names
That is my fault for not running the activateDAQ.py script after a round of rebuilds. I have run the script this morning, and confirmed that the oplev channels are showing up in dataviewer.

Quote:

Although Joe and Kiwamu claim that they have inserted the correct DAQ names for the OPLEVs (e.g. PERROR and YERROR) back in Jan. 11, when I look today, I see that these channels are missing!

I want my PERROR/YERRORs back!

 

 

  4299   Tue Feb 15 11:53:02 2011 kiwamuSummaryGeneraltoday's missions

I think this week is going to be an "alignment week".

The goal is to get a good alignment on X arm for both the green and the IR beam in order to reduce a2l couplings.

 

Today's missions are :

 - fixing the oplev channel names (see here)

 - fixing the oplev gain issue (see here)

 - engage the oplev servos

 - f2p adjustment

 - make a realtime lockin model for the f2p measurement and the dithering technique

 - alignment of the MC incident beam  (because we installed a new PMC this morning)

 - manual alignment of the IR beam by steering PZT1 and PZT2 (this procedure will be replaced by an automatic way soon)

 - bounce roll filters (see here)

 

  4298   Tue Feb 15 11:43:53 2011 JenneUpdateComputersOccasional error with NDS2

Just in case anyone has encountered this / knows how to fix it....

I'm running NDS2 on Rossa, trying to get a bunch of raw data from S5.  I get 10min of data at a time, and it goes through ~200 iterations successfully, and then throws the following error:

Getting new data
Connecting.... authenticate ... done
daq_recv_id: Wrong length read (0)
Error reading writerID in daq_recv_block
Warning: daq_request_data failed
 
??? Error using ==> NDS2_GetData
Fatal Error getting channel data.

Error in ==> getDARMdataTS at 37
oot = NDS2_GetData({...

Error in ==> SaveRawData_H1_DARM at 40
    oot = getDARMdataTS(t0s(ii), strideDuration, srate);

  4297   Tue Feb 15 06:59:57 2011 Suresh, JenneOmnistructureGeneralX end enclosure left open

[Jenne, Suresh]

 Jenne found the X-end table enclosure had been left open.  She replaced the lid on it.

 

  4296   Tue Feb 15 06:15:07 2011 SureshUpdateASCno signal from IP_ANG_Seg1

[Valery, kiwamu, Jenne, Suresh]

    I first interchanged the two QPD's on the Y end table to see if the problem QPD related.  Exchanging the units did not make any difference.  The problem therefore had to be in the cables or the circuit boards in 1X4

    We traced the signals pertaining to the IP_ANG QPD ( "Initial Pointing Beam") using  Jay's wiring diagram (pages 2 and 5 of 7).  We noted that while the signals were available on all Segments till the Monitors (Lemo) on 1X4-2-2A card, two of the lines did not reach the output of the cross connect 1X4-B8.  We checked card to make sure that the signals were indeed reaching the back plane of the 1X4-2 chassis using a D990612 extension board.  The card was found to be okay.  We therefore suspected that the cable (CAB_1X4_?) going from the card to the cross connect 1X4-B8 was faulty.  Indeed visual inspection showed that the crimping of the connector was poor and weight of the cable had put further strain on the crimping.  

   I changed the 64-pin connector on the 1X3-2-2A side of the cable. 

When I connected everything back together the problems persisted. Namely the lines P1-1A  (Segment 1 high) and P1-2C (Segment 2 Low) were floating They were not reaching points 2T and 3T respectively on the output of the cross connect.

   I therefore replaced 1X4-B8 with a similar unit which I found in one of the shelves along the East (Y) arm. 

I then checked with the StripTool to make sure that all the quadrants are showing similar response to a flashlight on the QPD.   All Segments are working fine now. Currently the IR Initial Pointing beam reaches the QPD but is not centered on it. 

I did not attempt to center it since the beam appeared to be clipped and may anyway require repositioning.

JD: We need to meditate on where this beam could be getting clipped.  Suresh and I checked that it's not on the viewport on the beam's way out of the ETMY chamber by seeing that the beam is far away from the edges of the viewport, and also far away from the edges of the black beamtube between the viewport and the table.  Suresh mentioned that the clipping nature of the IP_ANG beam sometimes goes away.  I don't know if this is the same clipping that Kiwamu might be seeing with the main beam, or if this is separate clipping just with the IP beam, after it's been picked off.  I suspect it's the same as what Kiwamu is seeing....maybe when we move PZT1, we clip on one of the MMT mirrors or PZT2??  If this is true, it's a total pain since we might have to vent if we can't steer around it.

Y-end_beam.JPG

 

  4295   Tue Feb 15 03:10:37 2011 kiwamuUpdateASCIR beam alignment for Xarm : TRX reduction

I tried aligning the IR beam axis for the X arm to have good beam centering on ITMX and ETMX.

As a first attempt, I started translating the beam upward by steering PZT1 and PZT2, since the pitch was quite off from the center on ITMX.

As a result I could decrease the pitch off-centering down to about 0.5 mm on ITMY, but on the other hand TRX decreased a lot (by a factor of 4).

I am worrying if something in the central part of IFO might be clipping the beam.

 


(notes)

When I was touching PZT1 and PZT2, I payed attention on IP_ANG so that I don't lose a beam spot on IP_ANG.

As long as the beam is on the IP_ANG QPD, the angle of the beam should not be so much different.

Each time after I touched the PZTs, I realigned ITMX and ETMX to maximize the transmitted light.

In this way I proceeded the alignment by changing the PZT offsets little by little while keeping the X arm locked always.

At the beginning, all the PZT offsets were zero. And at the end of this work they became:

 C1:LSC-PZT1_Y = 1.880

 C1:LSC-PZT2_Y = -1.699

But during this alignment work TRX gradually decreased eventually down to 0.25, which had been 1 at the beginning (TRX is calibrated by dividing it by its maximum power).

Along with this TRX reduction, I found that the optical gain also decreased by a factor of about 5.

This fact has been confirmed by intentionally increasing the filter gain such that the servo oscillates at the UGF.
 

Quote:

The amounts of the X arm's beam off-centering have been measured by the A2L technique.

     - ETMX

         PIT  = -1.61 mm

         YAW =  -0.918 mm

    - ITMX

         PIT = -3.76 mm

        YAW = -2.24 mm

 

  4294   Tue Feb 15 02:13:16 2011 kiwamuSummaryASCa daytime task : small signals on ETMX OL

Rana and I found that the QPD for the optical lever at X end are showing small signals.

At this moment each of the segments exhibits approximately 200 counts when the oplev beam is centered.

These small numbers may be due to the coating of ETMX, but we are not sure.

Probably we have to increase the gain of the QPD depending on situations.

 

So a set of the tomorrow's daytime task is:

   1. check the trend data of the QPD outputs to see how much signals were there in the past.

   2. check the whitening filters to make sure if it's on or off.

   3. If it's necessary, increase the gain of the QPD to have reasonable readouts.

I am going to ask somebody to do this task.

  4293   Mon Feb 14 23:29:04 2011 ranaSummarySUSETMX Fitlers moved around

All of the SUS used to have only 1 filter module for SIDE. They now have 3 filter modules for SIDE just like the other DOFs.

Today I moved the filters around so that the sensor filters are in SDSEN, the servo filters are in SUSSIDE, and the dewhitening for the coil is in SDCOIL.

I noticed along the way that the bounce/roll mode notches for all of the suspensions are still set for the frequencies of the previous suspensions. Suresh has 'volunteered' to find the new frequencies and make the new bandstop filters by looking up the seminal work on this by Dan Busby / Sam Waldman.

  4292   Mon Feb 14 21:59:35 2011 ranaUpdateCDSUpdated some DAQ channel names

Although Joe and Kiwamu claim that they have inserted the correct DAQ names for the OPLEVs (e.g. PERROR and YERROR) back in Jan. 11, when I look today, I see that these channels are missing!

I want my PERROR/YERRORs back!

 

  4291   Mon Feb 14 18:27:39 2011 josephbUpdateCDSBegan updating to latest CDS svn, reverted to previous state

[Joe, Alex]

This morning I began the process of bringing our copy of the CDS code up to date to the version installed at Livingston. The motivation was to get fixes to various parts, among others such as the oscillator part.   This would mean cleaning up front end model .mdl files without having to pass clk, sin, cos channels for every optic through 3 layers of simulink boxes.

I also began the process of using a similar startup method, which involved creating /etc/init.d/ start and stop scripts for the various processes which get run on the front ends, including awgtpman and mx_streams.  This allows the monitor software called monit to remotely restart those processes or provide a web page with a real time status of those processes.  A cleaner rc.local file utilizing sub-scripts was also adapted.

I did some testing of the new codes on c1iscey.  This testing showed a problem with the timing part of the code, with cycles going very long.  We think it has something to do with the code not accounting for the fact that we do not have IRIG-B timing cards in the IO chassis providing GPS time, which the sites do have.  We rely on the computer clock and ntpd.

At the moment, we've reverted to svn revision 2174 of the CDS code, and I've put the previously working version of the c1scy and c1x05 (running on the c1iscey computer) back. Its from the /opt/rtcds/caltech/c1/target/c1x05/c1x05_11014_163146 directory.  I've put the old rc.local file back in /diskless/root/etc/ directory on the fb machine.  Currently running code on the other front end computers was not touched.

  4290   Mon Feb 14 17:16:47 2011 steveUpdateGeneralracks, breakers labeled

The south arm labels were changed from y to x to reflect reality.

So racks, manual disconnects and breakers now have their actual name. The east arm will be changed over tomorrow.

Please remove incorrect labels if you see any !

I can not find 1X4 breaker so it will be  traced.

  4289   Mon Feb 14 15:59:49 2011 JenneUpdateIOOWFS quantum efficiency as a function of angle

[Larisa and Jenne]

A few weeks ago (on the 28th of January) I had tried to measure the quantum efficiency of one quadrant of the WFS as a function of angle.  However, Rana pointed out that I was a spaz, and had forgotten to put a lens in front of the laser.  Why I forgot when doing the measurement as a function of angle, but I had remembered while doing it at normal incidence for all of the quadrants, who knows?

Anyhow, Larisa measured the quantum efficiency today.  She used WFS2, quadrant 1 (totally oil-free), since that was easier than WFS1.  She also used the Jenne Laser (with a lens), since it's more stable and less crappy than the CrystaLasers.  We put a 50 Ohm terminator on the RF input of the Jenne Laser, since we weren't doing a swept sine measurement.  Again, the Ophir power meter was used to measure the power incident on the diode, and the reflected power, and the difference between them was used as the power absorbed by the diode for the quantum efficiency measurement.  A voltmeter was used to measure the output of the diode, and then converted to current as in the quote below. 

Still on the to-do list:  Replace the WFS2 diode.  See if we have one around, otherwise order one.  Align beams onto WFS so we can turn on the servo.

QE = (h*c)/(lambda*e) * (I/P)

Where I = (Volts from Pin1 to GND)/2 /500ohms
P = Power from laser - power reflected from diode.
h, c, e are the natural constants, and lambda is 1064nm.
Also, I/P = Responsivity


Larissa is going to put her data and plots into the elog shortly....

Quote:

Quantum Efficiency Measurement:

I refer to Jamie's LHO elog for the equation governing quantum efficiency of photodiodes: LHO 2 Sept 2009

The information I gathered for each quadrant of each WFS was: [1] Power of light incident on PD (measured with the Ophir power meter), [2] Power of light reflected off the PD (since this light doesn't get absorbed, it's not part of the QE), and [3] the photo current output by the PD (To get this, I measured the voltage out of the DC path that is meant to go to EPICS, and backed out what the current is, based on the schematic, attached). 

I found a nifty 25 pin Dsub breakout board, that you can put in like a cable extension, and you can use clip doodles to look at any of the pins on the cable.  Since this was a PD activity, and I didn't want to die from the 100V bias, I covered all of the pins I wasn't going to use with electrical tape.  After turning down the 100V Kepco that supplies the WFS bias, I stuck the breakout board in the WFS.  Since I was able to measure the voltage at the output of the DC path, if you look at the schematic, I needed to divide this by 2 (to undo the 2nd op amp's gain of 2), and then convert to current using the 499 Ohm resistor, R66 in the 1st DC path.  

I did all 4 quadrants of WFS1 using a 532nm laser pointer, just to make sure that I had my measurement procedure under control, since silicon PDs are nice and sensitive to green.  I got an average QE of ~65% for green, which is not too far off the spec of 70% that Suresh found.

I then did all 8 WFS quadrants using the 1064nm CrystaLaser #2, and got an average QE of ~62% for 1064 (58% if I exclude 2 of the quadrants....see below).  Statistics, and whatever else is needed can wait for tomorrow.

Problem with 2 quadrants of WFS2?

While doing all of this, I noticed that quadrants 3 and 4 of WFS2 seem to be different than all the rest.  You can see this on the MEDM screens in that all 6 other quadrants, when there is no light, read about -0.2, whereas the 2 funny quadrants read positive values.  This might be okay, because they both respond to light, in some kind of proportion to the amount of light on them.  I ended up getting QE of ~72% for both of these quadrants, which doesn't make a whole lot of sense since the spec for green is 70%, and silicon is supposed to be less good for infrared than green.  Anyhow, we'll have to meditate on this.  We should also see if we have a trend, to check how long they have been funny.

 

ELOG V3.1.3-